a storage battery contains a solution of h2so4 38%by weight. At the concentration vant hoff factor os2.50.At what temperature will the battery content freeze?

Grade:12

1 Answers

Vikas TU
6 years ago
 
From Freezing point formulae,
Tf = iKfm
Tf = 2.50*1.86*m
where m = molality of the soln.
Now 38 g of H2SO4 is preent in 100 g of H2SO4 soln.
Mass of solvent = 100 – 38 => 62 gms.
m = (38/98)/(62/1000) => 38000/(62*98) = 6.25
 
Hence, Tf = 2.50*1.86*6.25 = 29.0625 ‘CFrom Freezing point formulae,
Tf = iKfm
Tf = 2.50*1.86*m
where m = molality of the soln.
Now 38 g of H2SO4 is preent in 100 g of H2SO4 soln.
Mass of solvent = 100 – 38 => 62 gms.
m = (38/98)/(62/1000) => 38000/(62*98) = 6.25
 
Hence, Tf = 2.50*1.86*6.25 = 29.0625 ‘C